The linear regression equation for a data set is ˆy = − 4.1 + 1.6x. The actual value at  x = 4 is 5.6. What is the residual value at x=4?


residual:


1
Expert's answer
2021-11-23T15:21:28-0500

The difference between the observed data value and the predicted value (the value on the straight line) is the error or residual.


"y(4)=-4.1+1.6(4)=2.3"

"residual = observed \u2013 predicted"

"=5.6-2.3=3.3"

The residual value at "x=4" is "3.3."